Output (units) Total cost (R)

0 100

1 110

2 130

3 166

4 220

5 300

The marginal cost of the third unit is  

 



1
Expert's answer
2021-08-11T09:17:46-0400

The marginal cost of the third unit is:

"MC(3) = \frac{TC3 - TC2} {Q3-Q2} = \frac{166-130} {3-2} = 36."